OpenStudy (anonymous):

solve the system using elimination 2x-y+z=10 4x+2y-3z=10 x-3y+2z=8

OpenStudy (anonymous):

@tkhunny @SolomonZelman

OpenStudy (mrnood):

take the 3rd equation and rewrite it as x=????

OpenStudy (anonymous):

x=3y-2z+8

OpenStudy (tkhunny):

The x's look good to me. 4x-2y+2z=20 -4x-2y+3z=-10 -4x+12y-8z=-32 Add the second to the first. Add the third to the first. Continue in this way.